Today was (and is still) devoted to optical/detector tests. Dan, Hannes, and Shaul are working hard on getting polarization rotation data across our focal plane.
In the meantime, things are breaking. Our GPS system has some sort of burned component inside and one of the receivers no longer works. And the analog channels of our so-called "general housekeeping" boards continue to be problematic.
My solution to our gondola balance issue, a set of garage door extension springs, seems to be working -- in fact, working a little too well! Before, our inner frame wanted to tip forward of its own accord; now, the springs are pulling up on the inner frame. We'll eventually find the sweet spot.
Because the gondola was being used today for optical tests, we couldn't do any more work on baffles. Jeff and I are planning on installing more baffles tomorrow, so we did some prep work for that today.
Not much more to report from here. No pictures today, but I have to remember to get a set of pictures of the garage door springs.